OpenStudy (anonymous):

Determine if the Sum from infinity to n=1 of n^3/ 5^n

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ n^3 }{ 5^n}\]

OpenStudy (anonymous):

you want the sum? or you want to know if it converges?

OpenStudy (anonymous):

the sum if it convereges

OpenStudy (anonymous):

actually just if it converges

OpenStudy (anonymous):

ratio test will work nicely for this one i think

OpenStudy (anonymous):

when using the ration test i have the lim-> infinity \[\frac{ (n+1)^3 }{ 5n^3 }\]

OpenStudy (anonymous):

and I'm stuck

OpenStudy (anonymous):

numerator and denominator are both polynomials of degree 3

OpenStudy (anonymous):

No it's 5^n ?

OpenStudy (anonymous):

leading coefficient of the numerator is 1, leading coefficient of the denominator is 5 limit is \(\frac{1}{5}\)

OpenStudy (anonymous):

or with the ratio test yes.

OpenStudy (anonymous):

in other words, pretty much from your eyeballs you see that \[\lim_{n\to \infty}\frac{ (n+1)^3 }{ 5n^3 }=\frac{1}{5}\]

OpenStudy (anonymous):

(n+1)^3 is not equal to n^3 + 1^3 ???

OpenStudy (anonymous):

no it is not

OpenStudy (anonymous):

but it is still a polynomial of degree 3 with leading coefficient 1

OpenStudy (anonymous):

in fact it is \(n^3+3n^2+3n+1\) but all the terms of degree lower than 3 don't matter when you take the limit

OpenStudy (anonymous):

oh so they would all go to 0 anywat

OpenStudy (anonymous):

accept n^3

OpenStudy (anonymous):

\[\lim{n\to \infty}\frac{n^3+3n^3+3n+1}{5n^3}=\frac{1}{5}\] right?

OpenStudy (anonymous):

yes but that means it would diverge because an does not equal 0. The answer is converge

OpenStudy (anonymous):

this from whatever pre-calc class you had where you knew that the horizontal asymptote of a rational function, where the degree of the numerator equals the degree of the denominator, is the ratio of the leading coefficients

OpenStudy (anonymous):

so your limit is \(\frac{1}{5}\)and since \(\frac{1}{5}<1\) the sum converges

OpenStudy (anonymous):

no you got two things confused

OpenStudy (anonymous):

the TERMS themselves must go to zero, although that is not enough for the sum to converge but the ratio test you have to check that the limit of the ratios is less than 1, not zero

OpenStudy (anonymous):

\[\lim\frac{a_{n+1}}{a_n}<1\]

OpenStudy (anonymous):

Ohhhhhhh ok thanks!

OpenStudy (anonymous):

yw